# Mixed Precision Iterative Refinement for Linear Inverse Problems

### v1

@@ 5.3 Examples and Results
Mean Error Min Error Max Error SD Error
(1,1,1) $1$ $-141.5e\text{-}$ 14 $-205.5e\text{-}$ 20 $-142.4e\text{-}$ 14 $-143.8e\text{-}$ 14
$5$ $-154.9e\text{-}$ 15 00 $-132.0e\text{-}$ 13 $-142.5e\text{-}$ 14
$10$ $-154.7e\text{-}$ 15 00 $-131.9e\text{-}$ 13 $-142.4e\text{-}$ 14
(2,1,1) $1$ $-61.5e\text{-}$ 6 $-148.8e\text{-}$ 14 $-54.7e\text{-}$ 5 $-66.4e\text{-}$ 6
$5$ $-84.1e\text{-}$ 8 $-141.7e\text{-}$ 14 $-77.2e\text{-}$ 7 $-71.2e\text{-}$ 7
$10$ $-84.1e\text{-}$ 8 $-141.7e\text{-}$ 14 $-77.2e\text{-}$ 7 $-71.2e\text{-}$ 7
(2,2,1) $1$ $-61.5e\text{-}$ 6 $-142.5e\text{-}$ 14 $-54.8e\text{-}$ 5 $-66.4e\text{-}$ 6
$5$ $-85.2e\text{-}$ 8 $-157.7e\text{-}$ 15 $-76.7e\text{-}$ 7 $-71.3e\text{-}$ 7
$10$ $-85.6e\text{-}$ 8 $-157.7e\text{-}$ 15 $-76.7e\text{-}$ 7 $-71.4e\text{-}$ 7
(3,2,1) $1$ $-21.2e\text{-}$ 2 $-95.0e\text{-}$ 9 $-12.9e\text{-}$ 1 $-24.0e\text{-}$ 2
$5$ $-46.5e\text{-}$ 4 $-121.1e\text{-}$ 12 $-22.9e\text{-}$ 2 $-33.6e\text{-}$ 3
$10$ $-46.5e\text{-}$ 4 $-121.1e\text{-}$ 12 $-22.9e\text{-}$ 2 $-33.6e\text{-}$ 3
(3,2,2) $1$ $-21.2e\text{-}$ 2 $-95.1e\text{-}$ 9 $-12.9e\text{-}$ 1 $-24.0e\text{-}$ 2
$5$ $-46.5e\text{-}$ 4 $-132.4e\text{-}$ 13 $-22.9e\text{-}$ 2 $-33.6e\text{-}$ 3
$10$ $-46.5e\text{-}$ 4 $-121.0e\text{-}$ 12 $-22.9e\text{-}$ 2 $-33.6e\text{-}$ 3
(3,3,2) $1$ $-21.2e\text{-}$ 2 $-95.0e\text{-}$ 9 $-12.9e\text{-}$ 1 $-24.1e\text{-}$ 2
$5$ $-46.6e\text{-}$ 4 $-105.0e\text{-}$ 10 $-22.7e\text{-}$ 2 $-33.3e\text{-}$ 3
$10$ $-46.8e\text{-}$ 4 $-104.4e\text{-}$ 10 $-22.7e\text{-}$ 2 $-33.3e\text{-}$ 3
(4,3,2) $1$ $01.3e$ 0 $-61.0e\text{-}$ 6 $13.6e$ 1 $05.0e$ 0
$5$ $-23.7e\text{-}$ 2 $-84.0e\text{-}$ 8 $-18.5e\text{-}$ 1 $-11.3e\text{-}$ 1
$10$ $-23.5e\text{-}$ 2 $-83.1e\text{-}$ 8 $-19.9e\text{-}$ 1 $-11.4e\text{-}$ 1
(4,3,3) $1$ $01.3e$ 0 $-79.6e\text{-}$ 7 $13.6e$ 1 $05.0e$ 0
$5$ $-23.8e\text{-}$ 2 $-84.0e\text{-}$ 8 $-18.9e\text{-}$ 1 $-11.4e\text{-}$ 1
$10$ $-23.6e\text{-}$ 2 $-83.2e\text{-}$ 8 $01.0e$ 0 $-11.4e\text{-}$ 1
(4,4,3) $1$ $01.4e$ 0 $-61.0e\text{-}$ 6 $13.6e$ 1 $05.0e$ 0
$5$ $-11.0e\text{-}$ 1 $-82.8e\text{-}$ 8 $04.4e$ 0 $-15.5e\text{-}$ 1
$10$ $-24.8e\text{-}$ 2 $-81.6e\text{-}$ 8 $01.3e$ 0 $-11.7e\text{-}$ 1
(4,4,4) $1$ $02.0e$ 0 $-78.6e\text{-}$ 7 $14.7e$ 1 $07.4e$ 0
$5$ $-11.3e\text{-}$ 1 $-81.0e\text{-}$ 8 $04.2e$ 0 $-15.3e\text{-}$ 1
$10$ $-11.4e\text{-}$ 1 $-91.5e\text{-}$ 9 $02.6e$ 0 $-13.9e\text{-}$ 1
@@
(Spectra - reconstruction) - We now compare MP-IR versus AIR on the Spectra problem. As our interests have focused on the filtering properties of MP-IR on the Tikhonov problem, we did not investigate methodologies for determining regularization parameters nor did we consider any stopping criterion. Both topics deserve their own dedicated studies in the context of mixed precision computing for inverse problems. In our numerical investigation we chose 3 magnitudes of regularization parameters: 1e-1, 1e-2, and 1e-3 as well as two noise levels: $0.5\%$ and $3\%$ to consider the broad behavior of both methods. For method comparison, we compute the relative reconstructive error (RRE) defined by
@@
where $x$ and $x^{(k)}$ denote the true solution and $k^{th}$ approximate solution determined by MP-IR or AIR, respectively.
Table 3 provides the RREs at selected iterations and precisions for MP-IR as well as AIR as a function of $\alpha^{2}$ and the noise level of the problem. For succinctness in the table, we did not provide the results for $\alpha^{2}=$1e-1 as AIR and all MP-IR precision variants did about the same in terms of minimum and final RREs amongst the $10$ iterations considered for both noise levels. Specifically, the RREs for both methods were $\approx 0.62$.
When $\alpha^{2}$ was decreased by an order of magnitude we observed that AIR attained larger errors than all MP-IR variants and was not stable as evident by a large final RRE value at iteration $10$. This general behavior was also observed when $\alpha^{2}=$1e-3 indicating that MP-IR would be a superior method for these choices of $\alpha^{2}$. We also observed that as the 3 precision choices for MP-IR were changed away from fp64 that it would take more iterations to reach the minimum RRE value. However, the relative errors still remained stable compared to AIR for the same number of iterations.
Table 3: Spectra example: relative errors for MP-IR and AIR methods as a function of the regularization parameter and noise level. The minimum error and its corresponding iteration number in parentheses as well as the final error for the maximum number of iterations allowed (10) is provided. MP-IR errors are provided for various mixed precision combinations.
@@
Min. RRE (It.) Fin. RRE Min. RRE (It.) Fin. RRE
1e-2 AIR $0.504\,\,(1)$ $46.9e$ 4 $0.517\,\,(1)$ $54.2e$ 5
MP-IR (1,1,1) $0.497\,\,(7)$ $0.497$ $0.500\,\,(1)$ $0.500$
(2,1,1) $0.497\,\,(1)$ $0.497$ $0.500\,\,(3)$ $0.500$
(2,2,1) $0.497\,\,(1)$ $0.497$ $0.500\,\,(1)$ $0.500$
(3,2,1) $0.496\,\,(1)$ $0.497$ $0.500\,\,(1)$ $0.500$
(3,2,2) $0.496\,\,(1)$ $0.497$ $0.500\,\,(1)$ $0.500$
(3,3,2) $0.496\,\,(1)$ $0.497$ $0.500\,\,(1)$ $0.500$
(4,3,2) $0.497\,\,(3)$ $0.497$ $0.500\,\,(2)$ $0.500$
(4,3,3) $0.496\,\,(3)$ $0.496$ $0.500\,\,(6)$ $0.500$
(4,4,3) $0.497\,\,(2)$ $0.497$ $0.501\,\,(7)$ $0.501$
(4,4,4) $0.496\,\,(10)$ $0.496$ $0.501\,\,(4)$ $0.501$
1e-3 AIR $0.406\,\,(1)$ $127.4e$ 12 $0.998\,\,(1)$ $134.4e$ 13
MP-IR (1,1,1) $0.347\,\,(2)$ $0.347$ $0.373\,\,(1)$ $0.373$
(2,1,1) $0.347\,\,(10)$ $0.347$ $0.373\,\,(3)$ $0.373$
(2,2,1) $0.347\,\,(2)$ $0.347$ $0.373\,\,(2)$ $0.373$
(3,2,1) $0.347\,\,(3)$ $0.347$ $0.373\,\,(3)$ $0.373$
(3,2,2) $0.347\,\,(3)$ $0.347$ $0.373\,\,(3)$ $0.373$
(3,3,2) $0.347\,\,(7)$ $0.347$ $0.373\,\,(8)$ $0.373$
(4,3,2) $0.346\,\,(2)$ $0.347$ $0.373\,\,(6)$ $0.373$
(4,3,3) $0.346\,\,(5)$ $0.347$ $0.373\,\,(6)$ $0.373$
(4,4,3) $0.347\,\,(5)$ $0.348$ $0.372\,\,(5)$ $0.372$
(4,4,4) $0.344\,\,(8)$ $0.350$ $0.371\,\,(4)$ $0.374$
(Image Deblurring - reconstruction) - For our final example we consider the 2D image reconstruction problem which comes from the software package IRTools [10]. Similarly to the reconstruction of Spectra, we compare the behavior of AIR against MP-IR for various precision choices. The Hubble image for this problem in Figure 3a contains $512\times 512$ pixels which corresponds to a blurring operator of size $512^{2}\times 512^{2}$ making both the application of chop and the storage of the full problem untenable. However, because the point spread function (Figure 3b) is rank-1, the blurring matrix $A$ can be decomposed into a Kronecker product, each part of which is of reasonable size to compute its SVD:
@@
Figure 3: Image deblurring example: (a) true Hubble image ($512\times 512$ pixels), (b) PSF ($20\times 20$ pixels), (c) blurred and $1\%$ noised image ($512\times 512$ pixels).
Similarly to the numerical results shown for the Spectra reconstruction example, Table 4 provides the RREs at selected iterations and precision combinations for MP-IR as well as AIR as a function of $\alpha^{2}$ and the noise level of the problem. Here we note that we considered different noise levels: $1\%$ and $3\%$ and regularization parameters: 1e0, 1e-1, and 1e-2 compared to the Spectra example, though again for the same reasons as before, we only show results for the two smaller regularization parameters. The same general behaviors observed for the Spectra example were also observed here. In Figure 4 we show the best reconstructions for MP-IR at precision (1,1,1) and (4,4,3) compared to AIR. We note that there are no meaningful visual differences between the images.
Table 4: Image deblurring example: relative errors for MP-IR and AIR methods as a function of the regularization parameter and noise level. The minimum error and its corresponding iteration number in parentheses as well as the final error for the maximum number of iterations allowed (10) is provided. MP-IR errors are provided for various mixed precision combinations.
@@
$1\%$ $3\%$
Min. RRE (It.) Fin. RRE Min. RRE (It.) Fin. RRE
1e-1 AIR $0.189\,\,(1)$ $102.7e$ 10 $0.190\,\,(1)$ $108.1e$ 10
MP-IR (1,1,1) $0.189\,\,(1)$ $0.189$ $0.190\,\,(1)$ $0.190$
(2,1,1) $0.189\,\,(3)$ $0.189$ $0.190\,\,(7)$ $0.190$
(2,2,1) $0.189\,\,(3)$ $0.189$ $0.190\,\,(2)$ $0.190$
(3,2,1) $0.189\,\,(3)$ $0.189$ $0.190\,\,(4)$ $0.190$
(3,2,2) $0.189\,\,(4)$ $0.189$ $0.190\,\,(7)$ $0.190$
(3,3,2) $0.189\,\,(3)$ $0.189$ $0.190\,\,(8)$ $0.190$
(4,3,2) $0.189\,\,(3)$ $0.189$ $0.190\,\,(3)$ $0.190$
(4,3,3) $0.189\,\,(10)$ $0.189$ $0.190\,\,(7)$ $0.190$
(4,4,3) $0.189\,\,(3)$ $0.189$ $0.190\,\,(9)$ $0.190$
(4,4,4) $0.189\,\,(4)$ $0.189$ $0.190\,\,(8)$ $0.190$
1e-2 AIR $0.182\,\,(1)$ $192.0e$ 19 $0.243\,\,(1)$ $195.9e$ 19
MP-IR (1,1,1) $0.174\,\,(3)$ $0.174$ $0.184\,\,(4)$ $0.184$
(2,1,1) $0.174\,\,(3)$ $0.174$ $0.184\,\,(10)$ $0.184$
(2,2,1) $0.174\,\,(3)$ $0.174$ $0.184\,\,(3)$ $0.184$
(3,2,1) $0.174\,\,(4)$ $0.174$ $0.184\,\,(6)$ $0.184$
(3,2,2) $0.174\,\,(9)$ $0.174$ $0.184\,\,(6)$ $0.184$
(3,3,2) $0.174\,\,(2)$ $0.174$ $0.184\,\,(10)$ $0.184$
(4,3,2) $0.174\,\,(8)$ $0.174$ $0.184\,\,(7)$ $0.184$
(4,3,3) $0.174\,\,(6)$ $0.174$ $0.184\,\,(5)$ $0.184$
(4,4,3) $0.174\,\,(8)$ $0.174$ $0.184\,\,(7)$ $0.184$
(4,4,4) $0.174\,\,(6)$ $0.174$ $0.184\,\,(7)$ $0.184$
Figure 4: Image deblurring example: (a) MP-IR with precisions $(1,1,1)$ at iteration $7$, (b) AIR at iteration $1$, (c) MP-IR with precisions $(4,4,3)$ at iteration $8$ all using $\alpha^{2}=$1e-2 and at their minimum RRE iteration according to Table 4.

### v2

@@ 5.3 Examples and Results
Mean Error Min Error Max Error Std. Dev.
(1,1,1) $1$ $-142.0e\text{-}$ 14 $-191.1e\text{-}$ 19 $-133.8e\text{-}$ 13 $-146.0e\text{-}$ 14
$5$ $-154.0e\text{-}$ 15 00 $-131.4e\text{-}$ 13 $-141.8e\text{-}$ 14
$10$ $-154.2e\text{-}$ 15 $-213.1e\text{-}$ 21 $-131.6e\text{-}$ 13 $-142.0e\text{-}$ 14
(2,1,1) $1$ $-52.1e\text{-}$ 5 $-111.3e\text{-}$ 11 $-43.8e\text{-}$ 4 $-56.7e\text{-}$ 5
$5$ $-62.7e\text{-}$ 6 $-142.3e\text{-}$ 14 $-55.2e\text{-}$ 5 $-67.8e\text{-}$ 6
$10$ $-62.7e\text{-}$ 6 $-142.3e\text{-}$ 14 $-55.2e\text{-}$ 5 $-67.8e\text{-}$ 6
(2,2,1) $1$ $-52.1e\text{-}$ 5 $-111.3e\text{-}$ 11 $-43.8e\text{-}$ 4 $-56.7e\text{-}$ 5
$5$ $-62.7e\text{-}$ 6 $-131.3e\text{-}$ 13 $-55.2e\text{-}$ 5 $-67.8e\text{-}$ 6
$10$ $-62.7e\text{-}$ 6 $-131.3e\text{-}$ 13 $-55.2e\text{-}$ 5 $-67.8e\text{-}$ 6
(2,2,2) $1$ $-52.0e\text{-}$ 5 $-129.8e\text{-}$ 12 $-43.8e\text{-}$ 4 $-56.4e\text{-}$ 5
$5$ $-62.6e\text{-}$ 6 $-131.6e\text{-}$ 13 $-55.2e\text{-}$ 5 $-67.8e\text{-}$ 6
$10$ $-62.8e\text{-}$ 6 $-146.6e\text{-}$ 14 $-55.3e\text{-}$ 5 $-68.0e\text{-}$ 6
(3,2,1) $1$ $-35.6e\text{-}$ 3 $-81.2e\text{-}$ 8 $-26.0e\text{-}$ 2 $-21.4e\text{-}$ 2
$5$ $-44.6e\text{-}$ 4 $-101.6e\text{-}$ 10 $-22.1e\text{-}$ 2 $-32.6e\text{-}$ 3
$10$ $-44.6e\text{-}$ 4 $-101.6e\text{-}$ 10 $-22.1e\text{-}$ 2 $-32.6e\text{-}$ 3
(3,2,2) $1$ $-35.6e\text{-}$ 3 $-81.2e\text{-}$ 8 $-26.0e\text{-}$ 2 $-21.4e\text{-}$ 2
$5$ $-44.6e\text{-}$ 4 $-101.6e\text{-}$ 10 $-22.1e\text{-}$ 2 $-32.6e\text{-}$ 3
$10$ $-44.6e\text{-}$ 4 $-101.6e\text{-}$ 10 $-22.1e\text{-}$ 2 $-32.6e\text{-}$ 3
(3,3,2) $1$ $-35.1e\text{-}$ 3 $-81.2e\text{-}$ 8 $-26.0e\text{-}$ 2 $-21.3e\text{-}$ 2
$5$ $-46.0e\text{-}$ 4 $-111.3e\text{-}$ 11 $-22.3e\text{-}$ 2 $-32.9e\text{-}$ 3
$10$ $-46.0e\text{-}$ 4 $-111.4e\text{-}$ 11 $-22.3e\text{-}$ 2 $-32.9e\text{-}$ 3
(3,3,1) $1$ $-35.1e\text{-}$ 3 $-81.2e\text{-}$ 8 $-26.0e\text{-}$ 2 $-21.3e\text{-}$ 2
$5$ $-46.0e\text{-}$ 4 $-111.4e\text{-}$ 11 $-22.3e\text{-}$ 2 $-32.9e\text{-}$ 3
$10$ $-46.0e\text{-}$ 4 $-111.4e\text{-}$ 11 $-22.3e\text{-}$ 2 $-32.9e\text{-}$ 3
(3,3,3) $1$ $-21.9e\text{-}$ 2 $-81.0e\text{-}$ 8 $-14.3e\text{-}$ 1 $-26.7e\text{-}$ 2
$5$ $-31.6e\text{-}$ 3 $-102.1e\text{-}$ 10 $-22.4e\text{-}$ 2 $-34.5e\text{-}$ 3
$10$ $-49.5e\text{-}$ 4 $-107.5e\text{-}$ 10 $-21.6e\text{-}$ 2 $-32.5e\text{-}$ 3
@@
(Spectra - reconstruction) - We now compare MP-IR versus AIR on the Spectra problem. As our interests have focused on the filtering properties of MP-IR on the Tikhonov problem, we did not investigate methodologies for determining regularization parameters nor did we consider any stopping criterion. Both topics deserve their own dedicated studies in the context of mixed precision computing for inverse problems. In our numerical investigation we chose 3 magnitudes of regularization parameters: 1e-2, 1e-3, and 1e-4 as well as two noise levels: $0.5\%$ and $3\%$ to consider the broad behavior of both methods. For method comparison, we utilize the relative reconstructive error (RRE) defined by
@@
where $x$ and $x^{(k)}$ denote the true solution and $k^{th}$ approximate solution determined by MP-IR or AIR, respectively. Using the RRE, we focus on the stability of the solutions by considering what we will refer to as the stable RRE (sRRE) which is computed as the mean of the 3rd through the 10th iterates, whose iterate numbers were chosen to reflect relatively stable solution behavior. We also consider the standard deviation values for the same choice of iterates. In our numerical studies, the precision variants of MP-IR did not require more than 10 iterations to converge to an RRE value that would deviate more than the standard deviation of the sRRE.
Table 3 provides the sRRE and standard deviation values at various precisions for MP-IR as well as AIR as a function of the regularization parameter and the noise level of the problem. For succinctness in the table, we did not provide the results for $\alpha^{2}=1e\text{-}2$ as the presented choices resulted in smaller sRRE values for MP-IR when considered for both noise levels. When AIR was investigated closely, we found that it was stable over the considered 10 iterations with a sRRE of $-16.26e\text{-}$ 1 and standard deviation of $-54.0e\text{-}$ 5 for a choice of $\alpha^{2}=1e\text{-}1$ and $3\%$ noise, however, this is a significantly worse error than was attainable by the MP-IR method shown in the table indicating a superior MP-IR performance for these choices of $\alpha^{2}$.
We observed that the sRRE values for precision combinations involving fp16 for the preconditioner (i.e., $\text{Pr}_{1}=3$) could experience larger sRRE variation - this is shown graphically in Figure 3 as well as in the standard deviation columns of Table 3. Specifically, we note that while the sRRE values for all precision combinations were fairly consistent across the same level of noise, regularization parameter, and precision combination, the standard deviation of the sRRE was found to be closer to the unit round-off of Pr1. This aligns with theoretical expectations given that this is the same precision that preconditioner is computed in. This variation can be observed directly in Figure 3a for combination (3, 3, 3)333(3,\,3,\,3)( 3 , 3 , 3 ), and additionally for combinations (3, 3, 1)331(3,\,3,\,1)( 3 , 3 , 1 ) and (3, 3, 2)332(3,\,3,\,2)( 3 , 3 , 2 ) in Figure 3b. From this experiment it appears that variability in RRE can be reduced by balancing the choice of the preconditioner in lower precision with higher working and residual precisions.
Table 3: Spectra example: stable relative reconstructive error (i.e., the average RRE for iterations $3$ $10$) and the standard deviation for MP-IR and AIR methods as a function of the regularization parameter for $0.5\%$ and $3\%$ noise. MP-IR errors are provided for various mixed precision combinations.
@@
sRRE Std. sRRE sRRE Std. sRRE
1e-3 AIR $112.0e$ 11 $115.8e$ 11 $121.1e$ 12 $123.4e$ 12
MP-IR (1,1,1) $-13.446e\text{-}$ 1 $-179.7e\text{-}$ 17 $-13.632e\text{-}$ 1 $-161.5e\text{-}$ 16
(2,1,1) $-13.446e\text{-}$ 1 $-123.6e\text{-}$ 12 $-13.632e\text{-}$ 1 $-125.2e\text{-}$ 12
(2,2,1) $-13.446e\text{-}$ 1 $-126.1e\text{-}$ 12 $-13.632e\text{-}$ 1 $-121.7e\text{-}$ 12
(2,2,2) $-13.446e\text{-}$ 1 $-86.8e\text{-}$ 8 $-13.632e\text{-}$ 1 $-87.4e\text{-}$ 8
(3,2,1) $-13.446e\text{-}$ 1 $-61.6e\text{-}$ 6 $-13.632e\text{-}$ 1 $-79.3e\text{-}$ 7
(3,2,2) $-13.446e\text{-}$ 1 $-61.5e\text{-}$ 6 $-13.632e\text{-}$ 1 $-79.5e\text{-}$ 7
(3,3,2) $-13.445e\text{-}$ 1 $-51.7e\text{-}$ 5 $-13.633e\text{-}$ 1 $-67.6e\text{-}$ 6
(3,3,1) $-13.445e\text{-}$ 1 $-51.9e\text{-}$ 5 $-13.633e\text{-}$ 1 $-69.7e\text{-}$ 6
(3,3,3) $-13.442e\text{-}$ 1 $-47.5e\text{-}$ 4 $-13.634e\text{-}$ 1 $-46.6e\text{-}$ 4
1e-4 AIR $196.2e$ 19 $201.8e$ 20 $203.7e$ 20 $211.1e$ 21
MP-IR (1,1,1) $-12.129e\text{-}$ 1 $-163.4e\text{-}$ 16 $-13.289e\text{-}$ 1 $-162.7e\text{-}$ 16
(2,1,1) $-12.129e\text{-}$ 1 $-115.0e\text{-}$ 11 $-13.289e\text{-}$ 1 $-113.1e\text{-}$ 11
(2,2,1) $-12.129e\text{-}$ 1 $-116.8e\text{-}$ 11 $-13.289e\text{-}$ 1 $-111.4e\text{-}$ 11
(2,2,2) $-12.129e\text{-}$ 1 $-71.4e\text{-}$ 7 $-13.289e\text{-}$ 1 $-71.5e\text{-}$ 7
(3,2,1) $-12.129e\text{-}$ 1 $-52.2e\text{-}$ 5 $-13.289e\text{-}$ 1 $-54.0e\text{-}$ 5
(3,2,2) $-12.129e\text{-}$ 1 $-52.3e\text{-}$ 5 $-13.289e\text{-}$ 1 $-54.0e\text{-}$ 5
(3,3,2) $-12.128e\text{-}$ 1 $-41.3e\text{-}$ 4 $-13.289e\text{-}$ 1 $-41.1e\text{-}$ 4
(3,3,1) $-12.129e\text{-}$ 1 $-41.5e\text{-}$ 4 $-13.289e\text{-}$ 1 $-41.9e\text{-}$ 4
(3,3,3) $-12.124e\text{-}$ 1 $-31.5e\text{-}$ 3 $-13.277e\text{-}$ 1 $-31.4e\text{-}$ 3
Figure 3: Spectra example: graph of RRE vs. iteration number for the MP-IR results for various precision combinations for $3\%$ noise and $\alpha^{2}=1e\text{-}3$. Pane (a) focuses on precision combination (3, 3, 3)333(3,\,3,\,3)( 3 , 3 , 3 ) capturing the large variation in the yellow region compared to the other precision combinations. Pane (b) focuses on the variability of precision combinations (3, 3, 1)331(3,\,3,\,1)( 3 , 3 , 1 ) and (3, 3, 2)332(3,\,3,\,2)( 3 , 3 , 2 ), denoted in the blue region which is smaller than the yellow region in pane (a).
(Image Deblurring - reconstruction) - For our final example we consider the 2D image reconstruction problem which comes from the software package IR Tools [11]. Similarly to the reconstruction of Spectra, we compare the behavior of AIR against MP-IR for various precision combinations. The Hubble image for this problem in Figure 4a contains $512\times 512$ pixels which corresponds to a blurring operator of size $256^{2}\times 256^{2}$ making both the application of chop and the storage of the full problem untenable. However, because the point spread function (Figure 4b) is rank-1, the blurring matrix $A$ can be decomposed into a Kronecker product, each part of which is of reasonable size to compute its SVD in fp64, fp32, and fp16:
@@
Figure 4: Image deblurring example: (a) true Hubble image ($512\times 512$ pixels), (b) PSF ($31\times 31$ pixels), (c) blurred and $1\%$ noised image ($256\times 256$ pixels).
Similarly to the numerical results shown for the Spectra reconstruction example, Table 4 provides the sRRE and standard deviation values at various precision combinations for MP-IR as well as AIR as a function of the regularization parameter and the noise level of the problem. Here, we note that we considered different noise levels: $1\%$ and $5\%$ and regularization parameters: $-11e\text{-}$ 1, $-21e\text{-}$ 2, and $-31e\text{-}$ 3 compared to the Spectra example, though again for the same reasons as before, we only show results for the two smaller regularization parameters. The same general behaviors observed for the Spectra example were also observed here with respect to observing larger sRRE standard deviations when utilizing a lower precision preconditioner.
We do note the observed rise in the sRRE value for the precision combination (3, 3, 3)333(3,\,3,\,3)( 3 , 3 , 3 ) involving Pr$=3$ 3 when the smaller $\alpha^{2}=1e\text{-}3$ was used. We attribute this behavior, in part, to the large variability experienced when using fp16 in the preconditioner (or equivalently, the solution basis) which can be seen through the standard deviation of the sRRE values in Table 4. From observing the effect that the choice of regularization parameter $-21e\text{-}$ 2 versus $-31e\text{-}$ 3 can make, it appears that this effect can be reduced with more regularization, but with some loss of accuracy. For conciseness, we do not show reconstructed images using AIR and MP-IR since the reconstructions amongst the latter for the various precision combinations are indiscernible and the former is meaningless given it’s RRE behavior.
Table 4: Image deblurring example: stable relative reconstructive error (i.e., the average RRE for iterations $3$ $10$) and the standard deviation for MP-IR and AIR methods as a function of the regularization parameter for $1\%$ and $5\%$ noise. MP-IR errors are provided for various mixed precision combinations.
@@
$1\%$ $5\%$
sRRE Std. sRRE sRRE Std. sRRE
1e-2 AIR $31.9e$ 3 $34.6e$ 3 $39.2e$ 3 $42.3e$ 4
MP-IR (1,1,1) $-12.620e\text{-}$ 1 $-174.5e\text{-}$ 17 $-12.628e\text{-}$ 1 $-174.5e\text{-}$ 17
(2,1,1) $-12.620e\text{-}$ 1 $-87.2e\text{-}$ 8 $-12.628e\text{-}$ 1 $-85.7e\text{-}$ 8
(2,2,1) $-12.620e\text{-}$ 1 $-86.7e\text{-}$ 8 $-12.628e\text{-}$ 1 $-85.2e\text{-}$ 8
(2,2,2) $-12.620e\text{-}$ 1 $-86.0e\text{-}$ 8 $-12.628e\text{-}$ 1 $-85.5e\text{-}$ 8
(3,2,1) $-12.621e\text{-}$ 1 $-42.7e\text{-}$ 4 $-12.629e\text{-}$ 1 $-42.7e\text{-}$ 4
(3,2,2) $-12.621e\text{-}$ 1 $-42.7e\text{-}$ 4 $-12.629e\text{-}$ 1 $-42.7e\text{-}$ 4
(3,3,2) $-12.629e\text{-}$ 1 $-32.7e\text{-}$ 3 $-12.636e\text{-}$ 1 $-32.6e\text{-}$ 3
(3,3,1) $-12.629e\text{-}$ 1 $-32.7e\text{-}$ 3 $-12.636e\text{-}$ 1 $-32.6e\text{-}$ 3
(3,3,3) $-12.630e\text{-}$ 1 $-33.2e\text{-}$ 3 $-12.639e\text{-}$ 1 $-33.3e\text{-}$ 3
1e-3 AIR $113.9e$ 11 $121.4e$ 12 $121.8e$ 12 $125.2e$ 12
MP-IR (1,1,1) $-12.468e\text{-}$ 1 $-173.3e\text{-}$ 17 $-12.561e\text{-}$ 1 $-162.0e\text{-}$ 16
(2,1,1) $-12.468e\text{-}$ 1 $-71.2e\text{-}$ 7 $-12.561e\text{-}$ 1 $-88.6e\text{-}$ 8
(2,2,1) $-12.468e\text{-}$ 1 $-82.9e\text{-}$ 8 $-12.561e\text{-}$ 1 $-86.1e\text{-}$ 8
(2,2,2) $-12.468e\text{-}$ 1 $-71.0e\text{-}$ 7 $-12.561e\text{-}$ 1 $-87.7e\text{-}$ 8
(3,2,1) $-12.547e\text{-}$ 1 $-22.5e\text{-}$ 2 $-12.638e\text{-}$ 1 $-22.4e\text{-}$ 2
(3,2,2) $-12.547e\text{-}$ 1 $-22.5e\text{-}$ 2 $-12.638e\text{-}$ 1 $-22.4e\text{-}$ 2
(3,3,2) $-12.925e\text{-}$ 1 $-11.4e\text{-}$ 1 $-13.031e\text{-}$ 1 $-11.5e\text{-}$ 1
(3,3,1) $-12.925e\text{-}$ 1 $-11.4e\text{-}$ 1 $-13.031e\text{-}$ 1 $-11.5e\text{-}$ 1
(3,3,3) $-13.001e\text{-}$ 1 $-11.7e\text{-}$ 1 $-13.098e\text{-}$ 1 $-11.7e\text{-}$ 1